1(a) Given the universal set U={1,2,3,…,9} and the sets
A={1,2,3,4,5}; B={4,5,6,7}; C={1,3,5,7,9}; D={2,4,6,8}.

Find
(i)B-A
(ii)(A-D)c
(iii)(A⋂C)\B

Question 2(a)
Given the functions f and g be defined by f(x) = 5x-4 over
8x-7
and g(x) =3x2-4. Find the formula defining the composition function gοf and fog.

Question 2 (b)
Consider the relation R= {(1,a); (2,d); (3,a); (3,b); (3,d)} and S= {(b,x); (b,z), (c,y); (d,z)}.Find S-1 o R-1


Answer by Edwin McCravy(20054)

1(a) Given the universal set U={1,2,3,…,9} and the sets
A={1,2,3,4,5}; B={4,5,6,7}; C={1,3,5,7,9}; D={2,4,6,8}.

Find
(i)B-A
That's the set of all elements of B that are not elements of B

From B = {4,5,6,7}, we take out any of these A={1,2,3,4,5} which
B contains.  A contains 4 and 5, which A also contains, so we 
remove 4 and 5 from B and we have {6,7}.

Therefore B-A = {6,7}


                                                          
(ii)(A-D)c

We first find what's in the parentheses, A-D

From A={1,2,3,4,5}, we take out any of these D={2,4,6,8} which
A contains.  D contains 2 and 4, which A also contains, so we 
remove 2 and 4 from A and we have {1,3,5}.

So A-D = {1,3,5}

So to find the complement of that set, (A-D)c,
we form the set of all elements in the universal set U={1,2,3,…,9}
which are not elements of A-D, which is {1,3,5}. So we remove
1,3 and 5 from the universal set U={1,2,3,…,9} and we have
{2,4,6,7,8,9}

So (A-D)c = {2,4,6,7,8,9}

We are only supposed to answer 2 questions per post.

Edwin